Ba(OH)2 was dissolved in pure water at 5 degree celsius until a saturated solution was obtained. the pH of this solution was found to be 12.25

a) what is the solubility of Ba(OH)2 in pure water at this temperature? answer in mol/L

b) what is the Ksp of the Ba(OH)2 at this temperature?
Answer #1

* PH Balolt) 2 = 12.25 PH e Polt=14 Polt= 14-12-25= 1.75 pol= -log (OH) 1.75= -log(OH) -1.75= log (oh) TOH] = 0.01778 = 1.778